[Brahms-l] Shift report 20040105 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Mon Jan 05 2004 - 16:00:43 EST
------------------------------------------------------------
Date: Mon Jan  5 16:00:42 2004, Shift: 08:00-16:00
Supervisor: Oana Ristea, with: 
Log book#15 pages 24 - 24

Shift Summary:
------------------------------------------------------------
1. Beam from 8:30 until 9:00 and from 10:30 until 12:45.

2.Run 9376-FFS@12 deg, BFS@8 deg, D1=D2=D3=D4=1/8B - 84K tr6;
           MRS@90 deg, D5=150 B

  run 9377-non TPC pedestal run;

  run 9378-FFS@12 deg, BFS@8 deg, D1=D2=D3=D4=1/8B -256K tr6;
           MRS@90 deg, D5=150 B

3. DB fixed the problem with trigger 3 (see e-log).
   TR-FS has slat 4, 5 ineficient.
   Still BFS platform can not be moved.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 5 16:01:16 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 05 2004 - 16:01:30 EST